OpenStudy (anonymous):

Find the values of k that make 2x2 + kx + 5 factorable. 7 or 11 2 or 5 1 or 10 1 or 5

OpenStudy (anonymous):

I was thinking A

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

A is right

OpenStudy (solomonzelman):

you plugged the options in, correct ?

OpenStudy (anonymous):

yes I did do you want to see my work?

OpenStudy (solomonzelman):

I believe you, and you got it correctly:)`

OpenStudy (anonymous):

thank you ! :)

OpenStudy (solomonzelman):

I was going to post a prove: \(\large\color{black}{ \displaystyle k^2-4(2)(5) }\) \(\large\color{black}{ \displaystyle k^2-4(10) }\) \(\large\color{black}{ \displaystyle k^2-40 }\)

OpenStudy (solomonzelman):

I am doing a discriminant test, and you know that for it to be factor-able, the discriminant has to be a perfect square (like 1, 4, 9, 16, 25 and on...). And certainly it can't be negative, Therefore \(\large\color{black}{ \displaystyle k^2\ge40 }\)

OpenStudy (solomonzelman):

And ergo, A is the correct answer, because in other options you have at least 1 value (for k) that doesn't satisfy the statement \(\large\color{black}{ \displaystyle k^2\ge40 }\).

OpenStudy (solomonzelman):

hope this logic makes sense...

OpenStudy (solomonzelman):

yw, tho

OpenStudy (anonymous):

oh ok I see how you got it the discriminant of a function ax^2 + bx + c is b^2 - 4ac so... ur prob 2x^2 + kx + 5 ax ^2 + bx + c the discriminant is.. b^2 - 4ac k^2 -40 and this value should be above 0 which means.. k^2 - 40 > 0 k^2> 0 and k^2 - 40 should be a perfect square.... the choices that satisfy this conditions r.. 7 , 11

OpenStudy (solomonzelman):

well, i didn't say it should be above zero, as (although 0 is not considered a perfect square, as regards to this discriminant test) it does only give you that the trinomial is a perfect square trinomial. \(\large\color{black}{ \displaystyle k^2\ge40 }\) and not \(\large\color{red}{ \displaystyle k^2>40 }\) BUT, you got the concept

OpenStudy (solomonzelman):

well, the bottom line is...

OpenStudy (solomonzelman):

\(\normalsize\color{royalblue}{ \rm Discriminant ~=~perfect~square }\) the trinomial can be factored into integer roots. (perfect square implies that discriminant is AT LEAST 0) \(\normalsize\color{royalblue}{ \rm Discriminant ~>~0 }\) some real number solutions \(\normalsize\color{royalblue}{ \rm Discriminant ~=~0 }\) trinomial is a perfect square \(\normalsize\color{royalblue}{ \rm Discriminant ~<~0 }\) some imaginary solutions